Carlberg Company has two manufacturing departments, assembly and painting. The assembly department started 11,400 units during November. The following production activity unit and cost information refers to the assembly department’s November production activities.

Assembly Department Units Percent of Direct
Materials
Percent of
Conversion
Beginning work in process 2,000 75 % 25 %
Units transferred out 9,000 100 % 100 %
Ending work in process 4,400 85 % 35 %
Beginning work in process
inventory—Assembly dept
$ 2,186 (consists of $1,688 for direct materials
and $498 for conversion)
Costs added during the month:
Direct materials $ 17,422
Conversion $ 20,582

Required:
Calculate the assembly department’s cost per equivalent unit of production for materials and for conversion for November. Use the weighted-average method.
Answer #1

Equivalent units = Units transferred out + (Ending work in process X % of completion)

Material = 9,000 + (4,400 X 85%) = 12,740

Conversion = 9,000 + (4,400 X 35%) = 10,540

Cost per equivalent unit = Total cost / Equivalent units

Material = ($1,688 + $17,422) / $12,740 = $1.50

Conversion = ($498 + $20,582) / 10,540 = $2